OpenStudy (anonymous):

A sports club has 84 members....

OpenStudy (anonymous):

attaching the rest of the question...

OpenStudy (anonymous):

Heres the answer choices

OpenStudy (anonymous):

@pitamar @ranga @undeadknight26

OpenStudy (anonymous):

will medal and fan!!!

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

@perl

OpenStudy (anonymous):

i think its A

OpenStudy (anonymous):

Sorry I was helping someone else. So we are told 84 members play baseball, 42 of them also play basketball. That means that (84-42)=42 play only baseball. We are now told that 25 play only basketball and 8 play none. In total we have (42 + 42 + 25 + 8) = 117. Now let's take the ratio each group takes of the total: $$ \text{both} = \frac{42}{117} \approx 0.36 \\ \text{baseball_only} = \frac{42}{117} \approx 0.36 \\ \text{basketball_only} = \frac{25}{117} \approx 0.21 \\ \text{none} = \frac{8}{117} \approx 0.07 $$Which matches the values on table A. The only thing that bothers me is the totals on the right. They seem pretty messed up because, for instance, not all the people play baseball which means that the total of baseball players cannot be 1. It seems that those values are messed on all tables. I couldn't find a single solution that the totals made sense to me... so I'd go with A too.